OpenStudy (anonymous):

∫γ(xz2y+1)ds, where γ(t) =(lnt, t^2, 2t) and tϵ[1,e] please help

OpenStudy (anonymous):

x(t)=lnt y(t)=t^2 z(t)= 2t x'(t)= 1/t y'(t)=2t z'(t)=2

OpenStudy (anonymous):

ds=\[\sqrt{\frac{ 1 }{ t^2 }+4t^2+4}dt \]

OpenStudy (anonymous):

\[\int\limits_{1}^{e} \frac{ lnt*2t }{ 2t^2+1 }\sqrt{\frac{ 1 }{ t^2 }+4t^2+4} dt\]

OpenStudy (anonymous):

and here i'am stuck. please help

OpenStudy (blockcolder):

The expression under the radical is a perfect square.

OpenStudy (anonymous):

\[\frac{ 2t^2 +1}{ t }?\]

OpenStudy (blockcolder):

yep

OpenStudy (anonymous):

thank you. i was writing 4 instead of 1 and i didn't noticed the perfect square. the answer is 2.

OpenStudy (blockcolder):

no problem